1.     In a psychological experiment a researcher has six different groups with the following number of respondents: 25,30,18,27,28,22. Find

i) Mean (1 mks)

ii) Variance (3 mks)

iii Standard deviation (1 mks)

 


1
Expert's answer
2021-03-03T06:04:09-0500

Mean "\mu=\frac{25+30+18+27+28+22}{6}=\frac{150}{6}=25."

Variance "\sigma^2=\frac{(25-25)^2+(30-25)^2+(18-25)^2+(27-25)^2+(28-25)^2+(22-25)^2}{6}=\frac{96}{6}=16."

Standard deviation "\sigma=\sqrt{16}=4."